# MaskRIS: Semantic Distortion-aware Data Augmentation for Referring Image Segmentation

### v1

@@ 4.3 Ablation Study
Table 4: Impact of the masking ratio on the RefCOCO validation set. When a word $w_{i}$ is selected based on the masking ratio, it has a $p_{m}$ probability of being masked with a [MASK] token, a $p_{r}$ probability of being replaced with a random word from the vocabulary, and a $p_{u}$ probability of remaining unchanged.
Impact of Masking Ratio. We investigate how masking ratios affect model performance, focusing on the balance between masking images and text. Tab. 4 provides insights into the optimal use of masking in model training. For text masking, we follow the masking ratio used by BERT [7]. Note that we refer to the BERT setting as it is, setting $p_{u}$ to 10% is equivalent to combining $p_{m}$ to 89% with $p_{r}$ to 11%. We observe a significant drop in performance as the text masking ratio increases, indicating that removing too much textual information interferes with model training. On the other hand, image masking shows an optimal improvement at a masking ratio of 75%, suggesting that a higher degree of image masking improves model performance. These observations highlight the balance of masking ratios across different modalities. Excessive text masking can negatively affect learning by causing the loss of crucial information, whereas increasing image masking up to a certain threshold can be beneficial.
5 Conclusion
Supplementary Material
[…]
Contents
Figure A: Various types of mask sampling strategies. We use patch-wise sampling [11] as our default. For block-wise sampling, we follow BEiT[2] to remove large random blocks. For Cutout [8], we follow the implementation of timm [51]. For all strategies except Cutout, we maintain a consistent masking ratio of 75%.
Patch-wise Grid-wise [11] Block-wise [2] Cutout [8]
oIoU 76.49 76.14 75.90 76.07
Table A: Impact of mask sampling strategy. The performance (oIoU) is evaluated across various mask sampling strategies on the RefCOCO validation set.
Appendix A Ablation Study
A.1 Analysis of Mask Sampling Strategy
To rigorously validate the impact of image masking on model training, we explored various mask sampling strategies and their effects. Fig. A visually illustrates each method. Our baseline, patch-wise sampling [13, 15, 11] divides images into non-overlapping patches and randomly masks certain patches. This approach is widely used due to its simplicity and effectiveness. We also tested grid-wise sampling [11], where rows and columns are masked systematically rather than randomly. Additionally, we investigated block-wise masking, as used in BEiT [2], where contiguous regions of the image are masked, potentially eliminating more important structural details by covering larger areas. Finally, we evaluated Cutout [8], which masks a randomly selected rectangular section of the image.
Tab. A summarizes the results (oIoU) of each mask sampling strategy on the RefCOCO validation set. Remarkably, all the strategies we tested consistently outperform the previous SoTA method, CARIS [33] (74.67 oIoU), with significant margins. Among these strategies, our patch-wise random sampling approach achieved superior performance, demonstrating its ability to generate more varied training data than grid-wise sampling and minimize training data distortion more effectively than block-wise sampling. These results confirm the importance of selecting an appropriate masking strategy to improve performance.
Figure B: Training loss analysis of MaskRIS on RefCOCO. In the CARIS w/ Masking setting, we employ both image and text masking as data augmentation strategies within the CARIS [33] framework. We visualize (a) the training loss for the original (i.e., unmasked) inputs, (b) the training loss for the masked inputs, and (c) the performance of the model on the RefCOCO validation set.
@@ A.1 Analysis of Mask Sampling Strategy
CARIS [33] 50 74.67 77.63 71.71
CARIS [33] 100 74.80 77.38 70.63
@@
Table B: Comparison of computational cost. All results (oIoU) are evaluated on the RefCOCO dataset. MaskRIS outperforms CARIS even with half of the training epochs.
A.2 Analysis of Computational Cost
The MaskRIS framework introduces additional computational overhead due to its Distortion-aware Contextual Learning (DCL) mechanism, which processes both the original and masked inputs during training. To ensure a fair comparison, we adjusted the training epochs to evaluate whether the performance gains are simply due to increased computation. As shown in Tab. B, doubling the training epochs for CARIS does not improve its performance, confirming that MaskRIS’s superior results are not a attributable to increased computation. Furthermore, even when MaskRIS is trained for only 25 epochs, it still outperforms CARIS trained for 50 epochs. For computational efficiency, we measured the total training time on a setup with A6000$\times$2 GPUs. CARIS required 18.4 hours for 50 epochs, whereas MaskRIS achieved better performance in only 16.7 hours with 25 epochs. These results demonstrate the effectiveness of our masking strategy and DCL framework in achieving both high efficiency and strong performance. Through reducing the training time required for superior results, MaskRIS establishes itself as a practical and resource-efficient solution for improving model training.
A.3 Analysis of Training Loss
@@
We analyzed the MaskRIS’s effectiveness through an in-depth analysis of training loss, comparing three distinct settings: CARIS [33], CARIS w/ Masking, and MaskRIS. For CARIS w/ Masking, we employ both image and text masking as data augmentation strategies, applying them with a 50% probability. The training loss curves and performance trends over different epochs are shown in Fig. B. In the baseline CARIS, training loss with the original inputs rapidly converges, while training loss with the masked inputs increases over time. This indicates that CARIS tends to overfit to clean inputs as training progresses. Introducing masking in CARIS helps mitigate this overfitting by incorporating data augmentation. However, it disrupts learning stability and slows down loss convergence, limiting its effectiveness. In contrast, MaskRIS effectively resolves these issues by significantly improving loss convergence for both original and masked inputs while achieving superior performance. This is due to its ability to effectively mitigate the learning instability caused by input masking while maintaining the benefits of increased data diversity and robustness. These results highlight the robustness and efficiency of MaskRIS in optimizing training and improving model performance.
ps 8 16 32 64 112
oIoU 76.58 76.49 76.49 76.47 76.01
Table C: Impact of patch size on image masking. The performance (oIoU) is evaluated across different patch sizes on the RefCOCO validation set. ps denotes the patch size.
A.4 Impact of Hyperparameters
Patch Size of Image Masking. To further analyze the hyperparameters of image masking, we examined the training results of different patch sizes for patch-wise image masking. Tab. C shows that our method consistently achieves superior performance regardless of patch size, outperforming the current SoTA method, CARIS [33]. Specifically, we observe that a patch size of 8 yields the highest performance. Nevertheless, the results suggest that our method exhibits stable and strong performance as long as the patch size remains within a reasonable range, highlighting its robustness to patch size variations.
Loss Weight of DCL. As described in Sec. 3.3, our training objective of Distortion-aware Contextual Learning (DCL) is $\mathcal{L}_{total}=\mathcal{L}_{dist}+\mathcal{L}_{ce}.$ To balance the scale of the loss values with previous methods [33] during the training phase, we adapt the equation to:
$\mathcal{L}_{total}=\lambda\cdot\mathcal{L}_{dist}+(1-\lambda)\cdot\mathcal{L}_{ce},$ (5)
where $\lambda$ is set to 0.5 for all experiments to avoid extensive hyperparameter search for optimal performance. A $\lambda$ of 1.0 indicates training only with $\mathcal{L}_{ce}$, which is the equivalent to CARIS [33]. In Tab. D, we explore how varying $\lambda$ affects the DCL procedure. As shown in Tab. D, while the best performance is achieved with $\lambda=0.25$, the performance with $\lambda$ of 0.5 remains significantly better than the results without $\mathcal{L}_{dist}$. These results confirm that MaskRIS does not require extensive hyperparameter search for $\lambda$ and performs reasonably well with $\lambda$ of 0.5.
$\lambda$ 0.1 0.25 0.5 0.75 1.0
oIoU 75.48 76.62 76.49 76.16 74.67
Table D: Impact of $\lambda$ on the DCL procedure. The performance (oIoU) is evaluated across various $\lambda$ on the RefCOCO validation set.
@@ Appendix B More Qualitative Examples
Fig. C provides additional qualitative examples on the RefCOCO dataset across different occlusion scenarios as described in Sec. 3.2. We categorize our analysis into three cases: (a) intra-object context masking, where occlusions occur within the object of interest; (b) inter-object context masking, where occlusions occur within non-target objects; and (c) text context masking, where parts of textual descriptions are obscured.
@@
In scenario (a), CARIS [33] often struggles to recognize objects that are only partially visible, highlighting its difficulty in handling incomplete visual information. The challenge is not limited to occlusions of the target objects. In scenario (b), CARIS [33] also faces challenges when occlusions occur in non-target areas surrounding the target object, reducing its ability to identify the intended objects correctly. Similar limitations are observed in the text description domain, as shown in Fig. C(c). While CARIS can successfully recognize the target objects with complete text descriptions, its performance drops when only partial text information is available. In contrast, our MaskRIS demonstrates robustness across these occlusion scenarios by improving its comprehension and utilization of diverse contextual details in the image and text data.
@@
Fig. D shows more qualitative examples on the RefCOCO dataset. As shown in Fig. D(a), our MaskRIS accurately captures the target objects without over- or under-capturing them. In Fig. D(b), it demonstrates strong robustness to occlusion, effectively identifying partially visible objects. Furthermore, as highlighted in Fig. D(c), MaskRIS successfully aligns target objects with the given text descriptions. These results demonstrate that our approach significantly improves the robustness of the model and its ability to utilize textual clues for accurate object identification.
@@
We also provide failure cases in Fig. D(d). When the referred object in an image or the given text description is ambiguous or difficult to recognize, our approach sometimes struggles to identify the correct target. For instance, in the first example of the top row, MaskRIS fails to recognize the drawing (“flowers”) on the “bottle”, instead capturing a different object. Another failure occurs when text descriptions include contrasting relative positions, such as “left” and “right”, which confuse the model. Insufficient text descriptions also pose significant challenges. When text descriptions do not sufficiently characterize the target object such as “yellow” or “her” in the bottom row, MaskRIS captures the other objects as the target. These examples illustrate that while MaskRIS shows significant improvements in robustness and accuracy across various scenarios, it can still be limited in handling complex visual scenes or overly vague and ambiguous text descriptions.
@@
Figure C: Qualitative examples under various occluded contexts on the RefCOCO dataset. Although CARIS [33] tends to be inaccurate under occluded contexts, MaskRIS produces accurate predictions, demonstrating its robustness to occlusion and incomplete information. For text context masking, the word highlighted in red is masked.
@@
Figure D: More qualitative examples on the RefCOCO dataset. MaskRIS mitigates the limitations of CARIS [33] by (a) capturing target objects more precisely, (b) being robust to occlusions, and (c) effectively using various textual clues. (d) We also provide failure cases.
@@ Appendix C Implementation Details
Most of our experimental results are based on CARIS [33]. For the image encoder, we used the Swin-Base Transformer [35], pre-trained on ImageNet-22k [6], and for the text encoder, we employed BERT-Base [7]. The maximum length of the text is set to 20 words. We used the AdamW [36] optimizer with a weight decay of 0.01. We applied different learning rates of $51e-$ 5 and $41e-$ 4 to encoders and the others, respectively, with a polynomial learning rate schedule with a power of 0.9. The model was trained for 50 epochs with a batch size of 16, and the input images were resized to 448 $\times$ 448.
To validate the adaptability of MaskRIS with other RIS methods, we integrated it into several existing models in both fully and weakly supervised settings, as shown in Tab. 2. Importantly, we follow the original training recipes of each method without any changes to the original architecture. This simple integration demonstrates the flexibility of MaskRIS and its seamless compatibility with existing frameworks.

### v2

@@ A.5 Impact of Hyperparameters
Table G: Impact of the masking ratio on the RefCOCO validation set. When a word $w_{i}$ is selected based on the masking ratio, it has a $p_{m}$ probability of being masked with a [MASK] token, a $p_{r}$ probability of being replaced with a random word from the vocabulary, and a $p_{u}$ probability of remaining unchanged.
Impact of Masking Ratio. We investigate how masking ratios affect model performance, focusing on the balance between masking images and text. Table G provides insights into the optimal use of masking in model training. For text masking, we follow the masking ratio used by BERT (BERT). Note that we refer to the BERT setting as it is, setting $p_{u}$ to 10% is equivalent to combining $p_{m}$ to 89% with $p_{r}$ to 11%. We observe a significant drop in performance as the text masking ratio increases, indicating that removing too much textual information interferes with model training. On the other hand, image masking shows an optimal improvement at a masking ratio of 75%, suggesting that a higher degree of image masking improves model performance. These observations highlight the balance of masking ratios across different modalities. Excessive text masking can negatively affect learning by causing the loss of crucial information, whereas increasing image masking up to a certain threshold can be beneficial.
@@
CARIS (CARIS) 50 74.67 77.63 71.71
CARIS (CARIS) 100 74.80 77.38 70.63
@@
Table H: Comparison of computational cost. All results (oIoU) are evaluated on the RefCOCO dataset. MaskRIS outperforms CARIS even with half of the training epochs.
A.6 Analysis of Computational Cost.
MaskRIS needs additional computational costs due to the dual-path framework, which requires forwarding both the original and masked inputs. For a fair comparison, we adjust the training epochs and examine the impact on performance. As shown in Table H, doubling the training epochs for CARIS does not enhance its performance, indicating that MaskRIS’s improved performance is not a result of increased computation. Moreover, even when our method is trained for only 25 epochs, it still outperforms CARIS trained for 50 epochs. This highlights the efficacy of our masking strategy and dual-path approach in boosting performance efficiently. Notably, since MaskRIS does not change the model architecture or add additional parameters, it maintains the same inference time as the baseline.
Figure B: Training loss analysis of MaskRIS on RefCOCO. In the CARIS w/ Masking setting, we employ both image and text masking as data augmentation strategies within the CARIS (CARIS) framework. We visualize (a) the training loss for the original (i.e., unmasked) inputs, (b) the training loss for the masked inputs, and (c) the performance of the model on the RefCOCO validation set.
A.7 Analysis of Training Loss.
@@
We validate the effectiveness of MaskRIS through an in-depth analysis of training loss, comparing three distinct settings: CARIS (CARIS), CARIS w/ Masking, and MaskRIS. For CARIS w/ Masking, we employ both image and text masking as data augmentation strategies, applying them with a 50% probability. Figure B shows the training loss and the model performance changes over various training epochs. In the baseline CARIS, training loss with the original inputs rapidly converges, while training loss with the masked inputs increases. This indicates that CARIS overfits to clean inputs as training progresses. In contrast, CARIS w/ Masking effectively addresses this overfitting by incorporating input masking. However, it degrades loss convergence and disrupts learning stability. Our MaskRIS improves loss convergence for both original and masked inputs and at the same time significantly increases model performance.
@@ Appendix B More Qualitative Examples
Figure C provides additional qualitative examples on the RefCOCO dataset across different occlusion scenarios as described in Sec. 3.2. We categorize our analysis into three cases: (a) intra-object context masking, where occlusions occur within the object of interest; (b) inter-object context masking, where occlusions occur within non-target objects; and (c) text context masking, where parts of textual descriptions are obscured.
@@
In scenario (a), CARIS (CARIS) often struggles to recognize objects that are only partially visible, highlighting its difficulty in handling incomplete visual information. The challenge is not limited to occlusions of the target objects. In scenario (b), CARIS (CARIS) also faces challenges when occlusions occur in non-target areas surrounding the target object, reducing its ability to identify the intended objects correctly. Similar limitations are observed in the text description domain, as shown in Figure C(c). While CARIS can successfully recognize the target objects with complete text descriptions, its performance drops when only partial text information is available. In contrast, our MaskRIS demonstrates robustness across these occlusion scenarios by improving its comprehension and utilization of diverse contextual details in the image and text data.
@@
Figure D shows more qualitative examples on the RefCOCO dataset. As shown in Figure D(a), our MaskRIS accurately captures the target objects without over- or under-capturing them. In Figure D(b), it demonstrates strong robustness to occlusion, effectively identifying partially visible objects. Furthermore, as highlighted in Figure D(c), MaskRIS successfully aligns target objects with the given text descriptions. These results demonstrate that our approach significantly improves the robustness of the model and its ability to utilize textual clues for accurate object identification.
@@
We also provide failure cases in Figure D(d). When the referred object in an image or the given text description is ambiguous or difficult to recognize, our approach sometimes struggles to identify the correct target. For instance, in the first example of the top row, MaskRIS fails to recognize the drawing (“flowers”) on the “bottle”, instead capturing a different object. Another failure occurs when text descriptions include contrasting relative positions, such as “left” and “right”, which confuse the model. Insufficient text descriptions also pose significant challenges. When text descriptions do not sufficiently characterize the target object such as “yellow” or “her” in the bottom row, MaskRIS captures the other objects as the target. These examples illustrate that while MaskRIS shows significant improvements in robustness and accuracy across various scenarios, it can still be limited in handling complex visual scenes or overly vague and ambiguous text descriptions.
@@
Figure C: Qualitative examples under various occluded contexts on the RefCOCO dataset. Although CARIS (CARIS) tends to be inaccurate under occluded contexts, MaskRIS produces accurate predictions, demonstrating its robustness to occlusion and incomplete information. For text context masking, the word highlighted in red is masked.
@@
Figure D: More qualitative examples on the RefCOCO dataset. MaskRIS mitigates the limitations of CARIS (CARIS) by (a) capturing target objects more precisely, (b) being robust to occlusions, and (c) effectively using various textual clues. (d) We also provide failure cases.
@@ Appendix C Implementation Details
Most of our experimental results are based on CARIS (CARIS). For the image encoder, we used the Swin-Base Transformer (Swin), pre-trained on ImageNet-22k (imagenet), and for the text encoder, we employed BERT-Base (BERT). The maximum length of the text is set to 20 words. We used the AdamW (adamw) optimizer with a weight decay of 0.01. We applied different learning rates of 1e51e-5 and 1e41e-4 to encoders and the others, respectively, with a polynomial learning rate schedule with a power of 0.9. The model was trained for 50 epochs with a batch size of 16, and the input images were resized to 448 $\times$ 448.
To validate the adaptability of MaskRIS with other RIS methods, we integrated it into several existing models in both fully and weakly supervised settings, as shown in Table 2. Importantly, we follow the original training recipes of each method without any changes to the original architecture. This simple integration demonstrates the flexibility of MaskRIS and its seamless compatibility with existing frameworks.